OpenStudy (anonymous):

Hey just wanting to know how you can find a parabola given the focal and directrix

OpenStudy (anonymous):

Meant the equation of a parabola

OpenStudy (anonymous):

Sorry i don't know the question but i just wanted to say Battle Field Friends

OpenStudy (anonymous):

The vertex is half way between the Focal and the directrix

OpenStudy (anonymous):

\[y^2=4px\]

OpenStudy (anonymous):

Yes I know and so then I just plug in the numbers then?

OpenStudy (anonymous):

parabola takes the from of \( y^2 = 4ax \)

OpenStudy (anonymous):

So then how would I solve it?

OpenStudy (anonymous):

The distance between directix and focus is 2p

OpenStudy (anonymous):

I will step out and let saseal take over so you do not get confused.

OpenStudy (anonymous):

Because I need to know how to find the equation first

OpenStudy (anonymous):

Because I need to know how to find the equation first

OpenStudy (anonymous):

distance equation?

OpenStudy (anonymous):

Yes but is that it?

OpenStudy (anonymous):

Yes but is that it?

OpenStudy (anonymous):

you take the midpoint between focus and directix \[(\frac{ x _{2}-x _{1} }{ 2 })\] change it to y if your parabola opens the other side

OpenStudy (anonymous):

yup that distance is your p and you plug it into the formula

OpenStudy (anonymous):

Soo is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright thank you!

OpenStudy (anonymous):

welcome